The Ksp of CaF2 is 4x10^-11. What is the maximum concentration of Ca2+ possible in a 0.10 M NaF solution?


1
Expert's answer
2021-10-29T02:18:53-0400

Ksp = Ca2+ x (2F-)2


4.0x10-11 = Ca2+ x (2 x 0.1)2

4.0x10-11 = Ca2+ x 0.22

4.0x10-11 = Ca2+ x 0.04

Ca2+ = 4.0x10-11/0.04

= 1.0 x 10-9